# Plan Your Travel and Travel with Your Plan: Wide-Horizon Planning and Evaluation via LLM

### v1

@@ Abstract
Travel planning is a complex task requiring the integration of diverse real-world information and user preferences. While LLMs show promise, existing methods with long-horizon thinking struggle with handling multifaceted constraints and preferences in the context, leading to suboptimal itineraries. We formulate this as an $L^{3}$ planning problem, emphasizing long context, long instruction, and long output. To tackle this, we introduce Multiple Aspects of Planning (MAoP), enabling LLMs to conduct wide-horizon thinking to solve complex planning problems. Instead of direct planning, MAoP leverages the strategist to conduct pre-planning from various aspects and provide the planning blueprint for planning models, enabling strong inference-time scalability for better performance. In addition, current benchmarks overlook travel’s dynamic nature, where past events impact subsequent journeys, failing to reflect real-world feasibility. To address this, we propose Travel-Sim, an agent-based benchmark assessing plans via real-world travel simulation. This work advances LLM capabilities in complex planning and offers novel insights for evaluating sophisticated scenarios through agent-based simulation.
@@ 1 Introduction
Planning a trip xie2024travelplanner ; delarosa2024trippaltravelplanningguarantees is not a straightforward task even for humans, as it requires gathering up-to-date information from various sources, such as hotels, transportation, and points of interest (POI), and integrating this information with preferences and constraints to create a feasible itinerary. With the rapid development of Large Language Models (LLMs) brown2020languagemodelsfewshotlearners ; dubey2024llama ; achiam2023gpt , LLM agents zhu2023ghostminecraftgenerallycapable ; qin2023toolllmfacilitatinglargelanguage greatly improve their capabilities in planning, offering promising opportunities to address the intricate challenges of real-world travel planning.
Existing studies in travel planning chen2024travelagent ; tang2024synergizing ; hao2024largelanguagemodelssolve focus on simple commonsense constraints and coarse-grained user requirements, employing CoT reasoning to integrate collected information into actionable plans. However, they struggle to deliver compelling and viable plans due to their inability to effectively incorporate multifaceted real-world information and implicit constraints into a holistic planning process. For example, a travel plan that is feasible by satisfying the spatial constraints may turn into an unpleasant journey due to ignoring various other aspects, including the special preferences, stamina limitations, or repetitive routes.
Therefore, a well-crafted travel plan necessitates the incorporation of a broader range of information, which in turn requires LLMs to possess enhanced retrieval capabilities and the ability to perform complex planning. We formulate this kind of planning as an $L^{3}$ planning problem with long context (including multifaceted information), long instruction (including CoT and output formatting instruction), and long output (including planning process and a final result). Compared to logical problems, e.g., mathematical reasoning, that demand multi-step analytical processes with strict deductive coherence, $L^{3}$ planning problems require LLMs to simultaneously integrate multifaceted constraints and heterogeneous information, necessitating wide-horizon thinking from various aspects rather than long-horizon thinking. We conduct preliminary experiments to investigate the LLM zero-shot capability of solving the $L^{3}$ planning problem. We implement a naive wide-horizon thinking with aspect-aware question decomposition. Our results reveal that LLMs struggle with addressing $L^{3}$ planning problems with long-horizon thinking, compared to naive wide-horizon thinking.
@@
From the preliminary experiments, we also observe naive wide-horizon thinking exhibits limitations: 1) lack of inter-aspect associations; 2) dependence on artifactual guidance; 3) poor inference-time scalability. To address these limitations, we propose "Multiple Aspects of Planning" (MAoP). In MAoP, instead of directly planning, we introduce a strategist to conduct the pre-planning using a two-stage strategy. The strategist first analyzes the long context and decompose the planning request into various aspects to be considered. For each aspect, the strategist provides short guidance for the planning model to further analyze subsequently. In the second routing stage, the strategist integrates the independent aspects into a coherent planning blueprint, leading to better inference-time scaling performance with more aspects considered. In the actual planning of MAoP, the planning model concentrates on guidance of a single aspect per dialogue turn by following the pre-planning blueprint, progressively constructing a comprehensive planning process via multiple turns of dialogue.
@@
Figure 1: The comparison between long-horizon and wide-horizon thinking reveals distinct cognitive approaches. While long-horizon thinking involves deep exploration of a single reasoning trajectory, wide-horizon thinking incorporates heterogeneous information and constraints from long contexts by considering various aspects. It necessitates parallel consideration of multiple dimensions, which are subsequently integrated to generate comprehensive outputs.
@@
Although MAoP enhances the travel planning capability, the evaluation remains challenging: 1) travel is highly personalized and significantly influenced by real-world dynamics; 2) travel is a consistent process, where previous events can influence the remaining journey. Previous studies xie2024travelplanner ; chen2024travelagent ; shao2024chinatravelrealworldbenchmarklanguage introduce constraint pass rates as metrics, which only reflect the partial feasibility of the plan and the satisfaction of coarse-grained user requirements. These metrics ignore the real-world influence and consistency in real travel scenarios, poorly reflecting actual feasibility and attractiveness.
As the proof is in the pudding, we propose a novel agent-based evaluation framework to simulate a trip based on the real-world environment. A “traveler”, which can be either a group or an individual, functions as an LLM agent to travel by referring to the plan. We design an event-driven sandbox environment integrated with real-world information. For simple events like transiting to a new place, the traveler can leverage map API data to determine the optimal mode of transportation. To simulate complex events like sightseeing, we design an event agent to simulate the experience by integrating up-to-date travel blog posts. Besides considerations of the environment, we meticulously design the traveler agent to make it more aligned with real-world individuals. We design a variety of travelers with distinct features and a traveler stamina engine that simulates stamina exertion. Beyond the "objective" metrics (constraint pass rates), we introduce "subjective" metrics that the traveler evaluates the plan based on the travel experience. The traveler agents offer multi-granularity feedback to assess the experience at multiple levels, ensuring comprehensive and reliable evaluation.
@@
  • Plan Your Travel: We explore the dilemma of addressing the $L^{3}$ problem with long-horizon thinking and propose MAoP to boost the wide-horizon planning capability.
  • Travel with Your Plan: We propose an agent-based evaluation framework that simulates real-world travel to evaluate the feasibility and appeal of the travel plan, offering novel perspectives for assessing complex scenarios.
@@ 2 Preliminaries
As shown in Figure 1, previous studies wang2023planandsolvepromptingimprovingzeroshot ; yao2023treethoughtsdeliberateproblem ; yao2022react have primarily centered on developing CoT methods and their variants to solve mathematical problems requiring deep exploration along a single trajectory. Deepseek-R1 guo2025deepseekr1 further enhances the long-horizon reasoning ability by leveraging the Reinforcement Learning. Compared to mathematical problems that require long-horizon thinking, we formulate our travel planning problem as an $L^{3}$ planning problem, characterized by long context comprehension, long instruction processing, and long planning process. $L^{3}$ planning problem necessitates LLMs to: 1) extract pertinent information from long contexts (e.g., multiple tour guides, spatial information, user information, and more in Appendix 6); 2) conduct deliberate thinking over multi-dimensional constraints (e.g., real-world constraints and user preferences). $L^{3}$ planning problem does not require the model to reason deeply, as solving mathematical problems with CoTs, but necessitates considering multiple aspects simultaneously in a wide-horizon view. In this paper, we investigate the potential of conducting wide-horizon thinking to solve the $L^{3}$ planning problems.
@@
   w/ Artificial Guidance (Long) 23.3 36.2
   w/ Artificial Guidance (Wide) 31.9 44.1
   w/ Self-Gen. Guidance (Wide) 29.4 42.0
@@
   w/ Artificial Guidance (Long) 52.6 62.5
   w/ Artificial Guidance (Wide) 58.9 68.0
   w/ Self-Gen. Guidance (Wide) 55.2 64.1
@@ 2.2 Wide-Horizon Thinking with Aspect-Aware Guidance
To explore the influence of simultaneously considering various aspects, we conduct a preliminary experiment to examine the naive wide-horizon thinking in travel planning. We craft guidance (instruction) that includes a series of aspects to be considered when generating a travel plan. To conduct long-horizon thinking (zero-shot CoT) as the baseline, we combine the long context (user information, travel blog posts, spatial information, etc) with the artificial guidance as the input. For naive wide-horizon thinking, we decompose the original guidance into various aspects [^1] Different from the studies perez2020unsupervisedquestiondecompositionquestion ; min-etal-2019-multi that decompose the question into simpler sub-questions to solve it step by step, we decompose the question into aspects to be considered concurrently.. LLMs independently follow corresponding guidance to analyze from different aspects and integrate the whole analysis to produce the output. In Table 1, we find that naive wide-horizon thinking with aspect-aware guidance significantly improves the travel plan quality with better feasibility and personalization scores. Although the wide-horizon thinking demonstrates significant potential, well-designed artificial guidance can further enhance the performance over self-generated guidance. More details can be checked in Section 5 and Appendix D.
3 Plan Your Travel: Wide-Horizon Thinking on Real-world Planning
Based on the preliminaries, we identify three key deficiencies to be improved to enhance the naive wide-horizon thinking: 1) Although aspect-aware thinking is concurrent, its independence inherently limits the capture of inter-aspect associations; 2) the artificial guidance demonstrates superiority, yet its deliberate design remains challenging to implement; 3) as the number of aspects increases, the performance can not be scaled with more inference-time compute. Therefore, we propose the Multiple Aspects of Planning (MAoP), a method that leverages a strategist to conduct the pre-planning to address these deficiencies in naive wide-horizon thinking.
@@ 3.1 MAoP Training
The training of MAoP consists of three stages: reward model training, rejection sampling finetuning for the strategist model, and RL training for the planning model. As shown in Figure 2, we mainly focus on the last two stages and provide more details in Appendix E.
Rejection Sampling Finetuning (RFT) for Strategist Model
In MAoP, as shown in Figure 2, the strategist mainly does two things in pre-planning: 1) decomposing the original instruction and generating the aspect-aware guidance; 2) routing the planning trajectories to capture inter-aspect associations. These core operations solve the mentioned deficiencies in naive wide-horizon thinking. To train such a strategist model, we implement RFT while keeping the planning model frozen  [^2] We do not further implement RL on the strategist model. The RL pipeline has to go through a frozen planning model to get rewards, making it hard to optimize.. For each long context and question, we prompt the strategist model to first conduct pre-planning for $N$ times, and subsequently the planning model to generate final plans. The rejection strategy is to reject the trajectory where all $N$ plans receive reward scores below the threshold.
RL Training for Planning Model
Before RL training, similar to training the strategist model, we first train the planning model with RFT using the same rejection strategy. After RFT (cold start), we conduct online RL training on the planning model to further improve aspect-aware thinking ability. We use GRPO shao2024deepseekmathpushinglimitsmathematical with two kinds of outcome rewards. The first reward comes from the same reward model in the RFT, and the second one is an additional travel plan format reward. We use the KL penalty to prevent severe biases. For loss calculation, we mask the aspect-aware guidance from the strategists, preventing the planning models from generating the guidance by themselves.
3.2 MAoT Inference
@@ 3.2.1 Pre-Planning
As shown in Figure 2, the strategist model accepts the long context and question as the input and decomposes the planning request into various aspects. For each aspect, the strategist generates concise guidance to instruct the planning model how to conduct thorough analysis. By parallel sampling the strategist multiple times, we can derive a large amount of aspect-guidance pairs.
@@
Different from the naive wide-horizon thinking that treats the aspects independently, the strategist additionally selects and aggregates the aspects to route the best planning blueprint. As the experiments in Appendix D.2, we find that if we directly increase the number of aspects by sampling the strategist more times, the planning model only benefits from considering 3 ~5 aspects without further inference-time scaling capability. It is because more aspects introduce more information but also more noise. To address this, the strategist aggregates the number of aspects into a smaller number and constructs a coherent planning blueprint. As shown in Figure 2, subsequent aspect guidance is determined through the influence of preceding multiple aspects. The routing process shifts the burden of considering numerous aspects from the planning model to the strategist model, thus enabling the scalability of considering more aspects to further improve the wide-horizon thinking.
@@ 3.2.2 Aspect-Aware Thinking
Based on the planning blueprint constructed by the strategist, the planning model sequentially conducts thinking over the aspects in a way of coherent multi-turn dialogue. With aspect-specific guidance, the planning model can conduct a more focused and in-depth analysis over the long context from this aspect. After multiple turns of profound analysis, the planning model produces the final plan based on the previous wide-horizon thoughts in the last turn of the conversation.
@@ 3.3 MAoP Distillation for One-Step Wide-Horizon Thinking
The implementation of MAoP requires two different models and a complicated planning process in multiple turns of dialogue. To accelerate and simplify this process, we directly finetune the models with the MAoP planning synthetic data. Inspired by Deepseek-R1 guo2025deepseekr1 , we especially finetune (SFT) the smaller models like Qwen-2.5 3B yang2024qwen25 , Llama-3.2-Instruct 3B grattafiori2024llama3 , and Deepseek-R1-Distill Qwen-7B guo2025deepseekr1 using 2K synthetic MAoP samples to empower the wide-horizon thinking ability for smaller models. To be specific for data construction, we only preserve the aspect-specific guidance after the routed by the strategist. We compress the entire aspect-aware thinking and the final aggregation into a single turn of conversation as the model output. Therefore, during the inference, the distilled models implement one-step wide-horizon thinking to plan from multiple aspects.
4 Travel with Your Plan: Consistent Evaluation via Agent-based Simulation
Despite the detailed plans we have developed by MAoP, their feasibility in the real world and whether they meet the personalized needs require further verification. As shown in Table 2, previous benchmarks design rule-based/LLM-based metrics, such as pass rates of different constraints. However, travel is a consistent process where previous events influence the remaining journey. By neglecting this consistency, such benchmarks fail to evaluate the actual feasibility and personalized needs in real-world travel scenarios.
As the proof is in the pudding, we design a benchmark using agent-based simulation, dubbed Travel-Sim. In Travel-Sim, the traveler agent experiences a simulated travel referring to the plan and offers feedback from the experience. We use the advanced model, Gemini 2.5-Pro-Exp-0325 google2024gemini , to roleplay the traveler. This simulation mechanism naturally addresses the consistency issue and provides personalized feedback with multiple levels of granularity from the travelers’ perspective.

### v2

# Wide-Horizon Thinking and Simulation-Based Evaluation for Real-World LLM Planning with Multifaceted Constraints

@@ Abstract
Unlike reasoning, which often entails a deep sequence of deductive steps, complex real-world planning is characterized by the need to synthesize a broad spectrum of parallel and potentially conflicting information and constraints. For example, in travel planning scenarios, it requires the integration of diverse real-world information and user preferences. While LLMs show promise, existing methods with long-horizon thinking struggle with handling multifaceted constraints, leading to suboptimal solutions. Motivated by the challenges of real-world travel planning, this paper introduces the Multiple Aspects of Planning (MAoP), empowering LLMs with "wide-horizon thinking" to solve planning problems with multifaceted constraints. Instead of direct planning, MAoP leverages the strategist to conduct pre-planning from various aspects and provide the planning blueprint for planners, enabling strong inference-time scalability by scaling aspects to consider various constraints. In addition, existing benchmarks for multi-constraint planning are flawed because they assess constraints in isolation, ignoring causal dependencies within the constraints, e.g, travel planning, where past activities dictate future itinerary. To address this, we propose Travel-Sim, an agent-based benchmark assessing plans via real-world simulation, thereby inherently resolving these causal dependencies. This paper advances LLM capabilities in complex planning and offers novel insights for evaluating sophisticated scenarios through simulation.
@@ 1 Introduction
Large Language Models (LLMs) wang2023planandsolvepromptingimprovingzeroshot ; erdogan2025plan ; shridhar2021alfworldaligningtextembodied ; huang2024understandingplanningllmagents have shown significant promise in planning by generating action sequences to achieve specified goals. However, transitioning from controlled environments to the complexities of real-world planning presents a formidable challenge xie2024travelplanner ; chen2024travelagent ; huang2024understandingplanningllmagents , defined by the need to manage a multitude of diverse and simultaneous constraints. Prevailing methods yao2022react ; prasad2024adaptasneededdecompositionplanning often rely on task decomposition, a linear and sequential methodology that breaks a complex problem into simpler sub-tasks. This approach is effective in domains with limited constraints, such as in GUI automation nguyen2025guiagentssurvey ; wang2025guiagentsfoundationmodels , but its feasibility diminishes sharply in real-world scenarios where constraints are deeply interconnected. The failure of this step-by-step strategy highlights a fundamental mismatch between the problem-solving approach and the non-linear nature of the challenge itself.
The cognitive model underpinning this sequential approach is analogous to that of logical and mathematical reasoning shao2024deepseekmathpushinglimitsmathematical ; guan2025rstarmathsmallllmsmaster ; ahn2024largelanguagemodelsmathematical . These domains epitomize long-horizon thinking: a deep, step-by-step deductive process that the entire chain of reasoning funnels toward a limited set of deterministic outcomes. Real-world planning chen2024travelagent ; hao2025largelanguagemodelssolve ; zhang2024askbeforeplanproactivelanguageagents , however, demands a fundamentally different cognitive paradigm. It requires an LLM to move beyond linear deduction to instead simultaneously integrate multifaceted constraints and heterogeneous information, necessitating wide-horizon thinking from various aspects rather than long-horizon thinking. Success hinges less on the deductive soundness of a sequence and more on the holistic feasibility of a plan that concurrently satisfies a wide array of interconnected variables.
In this paper, we explore the limitations and potential of LLMs in real-world planning through the lens of the travel planning scenario. We conduct preliminary experiments to investigate the LLM’s zero-shot capability of solving the travel planning problem. Alongside long-horizon baselines, we implement a simple wide-horizon approach based on aspect-aware question decomposition. Our results indicate that LLMs struggle to address real-world planning problems using long-horizon thinking, performing notably better with even a naive wide-horizon method.
@@
Our preliminary experiments also reveal that this naive wide-horizon approach, which relies on decomposing the request into various aspects, has its own significant limitations: 1) lack of inter-aspect associations; 2) dependence on well-crafted artifactual guidance; 3) poor inference-time scalability. To address these limitations, we propose "Multiple Aspects of Planning" (MAoP). In MAoP, instead of directly planning, we introduce a strategist to conduct the pre-planning using a two-stage strategy. The strategist first analyzes the context and decomposes the planning request into various aspects to be considered. For each aspect, the strategist provides short guidance for the planner to further analyze subsequently. In the second routing stage, the strategist integrates the independent aspects into a coherent planning blueprint, leading to better inference-time scaling performance with more aspects considered. In the actual planning of MAoP, the planner concentrates on guidance of a single aspect per dialogue turn by following the pre-planning blueprint, progressively constructing a comprehensive planning process via multiple turns of dialogue.
@@
Figure 1: The comparison between long-horizon and wide-horizon thinking reveals distinct cognitive approaches. While long-horizon thinking involves deep exploration of a single reasoning trajectory, wide-horizon thinking incorporates heterogeneous information and constraints in long contexts by considering various aspects. It necessitates parallel consideration of multiple dimensions, which are subsequently integrated to generate comprehensive outputs.
@@
While MAoP improves real-world planning, the evaluation of multi-constraint planning remains a major hurdle. The challenges (e.g, from travel planning) are twofold: 1) The subjective nature of travel planning means there is no universal optimal solution, as users weigh constraints like cost and convenience differently, making objective evaluation elusive. 2) The causal dependency within a journey means a single failure can dynamically violate a cascade of subsequent, interconnected constraints, rendering simple, static evaluations inadequate. Previous studies xie2024travelplanner ; chen2024travelagent ; shao2024chinatravelrealworldbenchmarklanguage introduce constraint pass rates as metrics, which only reflect the partial feasibility of the plan and the satisfaction of coarse-grained user requirements. These metrics ignore the real-world influence and causal consistency in real travel scenarios, poorly reflecting actual feasibility.
As the proof is in the pudding, we propose a novel agent-based evaluation framework to simulate a trip based on the real-world environment. At its core, an LLM-powered traveler agent executes a plan within an event-driven sandbox. By leveraging live traffic data from maps and qualitative insights from travel blogs, we enable the simulation to organically capture the dynamics and unforeseen events of real-world scenarios. To better emulate real-world individuals, the traveler agent is meticulously designed with diverse personas and a dynamic stamina engine that simulates physical exertion. Beyond the "static" metrics (constraint pass rates), we introduce "dynamic" metrics that the traveler evaluates the plan based on their experience, offering multi-granularity feedback to assess the experience at multiple levels.
@@
  • We propose MAoP to enhance wide-horizon thinking capabilities for solving real-world planning with multifaceted constraints.
  • We propose a simulation-based evaluation framework to evaluate the feasibility and appeal of the travel plan, offering novel insights for assessing complex scenarios.
@@ 2 Preliminaries
As shown in Figure 1, previous studies wang2023planandsolvepromptingimprovingzeroshot ; yao2023treethoughtsdeliberateproblem ; yao2022react have primarily centered on developing CoT methods and their variants to solve planning and reasoning problems requiring deep exploration along a single trajectory. Deepseek-R1 guo2025deepseekr1 further enhances the long-horizon reasoning ability by leveraging Reinforcement Learning. In contrast, real-world planning such as travel planning necessitates LLMs to: 1) extract pertinent information from long contexts (e.g., tour guides, spatial information, traveler information, etc); 2) conduct deliberate thinking over multifaceted constraints (e.g., real-world constraints and preferences). Travel planning problem does not require the model to reason deeply, but necessitates considering multiple aspects simultaneously in a wide-horizon view. In this paper, we investigate the potential of conducting wide-horizon thinking compared to long-horizon thinking in real-world planning.
@@
   w/ Artificial Guidance + CoT 23.3 36.2
   w/ Artificial Guidance + Plan&Solve 25.0 39.7
   w/ Artificial Guidance + Wide 31.9 44.1
   w/ Self-Gen. Guidance + Wide 29.4 42.0
@@
   w/ Artificial Guidance + CoT 52.6 62.5
   w/ Artificial Guidance + Plan&Solve 57.4 64.2
   w/ Artificial Guidance + Wide 58.9 68.0
   w/ Self-Gen. Guidance + Wide 55.2 64.1
@@ 2.2 Wide-Horizon Thinking with Aspect-Aware Guidance
We conducted a preliminary experiment to compare the efficacy of long-horizon versus wide-horizon thinking on real-world planning with multifaceted constraints. The input provided to the model comprises two components: a rich context and a structured guidance prompt. The context includes real-world information such as traveler profiles, travel blogs, and spatial data (see Appendix 6 for processing details). The guidance is a carefully crafted instruction that outlines the key aspects to consider when generating a travel plan. To establish our long-horizon thinking baselines, we use this guidance to prompt two methods: zero-shot CoT and the Plan-and-Solve framework wang2023planandsolvepromptingimprovingzeroshot , which relies on task decomposition. Our approach to naive wide-horizon thinking is centered on aspect-based decomposition. Critically, unlike prior works perez2020unsupervisedquestiondecompositionquestion ; erdogan2025plan that decompose a question into a sequence of simpler sub-tasks, our method breaks the guidance into multiple aspects designed to be considered concurrently. The LLM then independently analyzes each aspect and finally synthesizes the insights from these parallel analyses to generate an integrated output. In Table 1, we find that naive wide-horizon thinking with aspect-aware guidance significantly improves the travel plan quality with better feasibility and personalization scores. Although the wide-horizon thinking demonstrates significant potential, well-designed artificial guidance can further enhance the performance over self-generated guidance. More details can be checked in Section 5 and Appendix D.
3 Wide-Horizon Thinking on Real-world Planning
Based on our preliminary experiments, we find that while the naive wide-horizon thinking approach shows promise, it still exhibits several key deficiencies: 1) Although aspect-aware thinking is concurrent, its independence inherently limits the capture of inter-aspect associations; 2) the artificial guidance demonstrates superiority, yet its deliberate design remains challenging to implement; 3) as the number of aspects increases, the performance can not be scaled with more inference-time compute. Therefore, we propose the Multiple Aspects of Planning (MAoP), a method that leverages a strategist to conduct the pre-planning to address these deficiencies in naive wide-horizon thinking.
@@ 3.1 MAoP Training
The training of MAoP consists of three stages: reward model training, rejection sampling finetuning for the strategist  [^1] We do not further implement RL on the strategist. The RL pipeline has to go through a frozen planner to get rewards, making it hard to optimize., and RL training for the planner. As shown in Figure 2, we mainly focus on the last two stages and provide more details in Appendix E.
Rejection Sampling Finetuning (RFT) for Strategist
In MAoP, as shown in Figure 2, the strategist mainly does two things in pre-planning: 1) decomposing the original instruction and generating the aspect-aware guidance; 2) routing the planning trajectories to capture inter-aspect associations. These core operations solve the mentioned deficiencies in naive wide-horizon thinking. To train such a strategist, we implement RFT while keeping the planner frozen. For each request, we prompt the strategist to first conduct pre-planning for $N$ times, and subsequently the planner to generate final plans. The rejection strategy is to reject the trajectory where all $N$ plans, as defined in Eq. 1, fall below a predefined threshold.
RL Training for Planner
Following an initial cold-start RFT, we conduct RL training using GRPO shao2024deepseekmathpushinglimitsmathematical on the planner to further improve aspect-aware thinking ability. To mitigate reward hacking, we design a multi-dimensional reward function. It is primarily guided by the PER score (a composite of five criteria, see Sec 5.1 and Appendix E) and also includes an auxiliary reward for proper plan formatting. The overall function is defined as follows:
[…]
3.2 MAoP Inference
@@ 3.2.1 Pre-Planning
As shown in Figure 2, the strategist accepts the long context and question as the input and decomposes the planning request into various aspects. For each aspect, the strategist generates concise guidance to instruct the planner how to conduct a thorough analysis. By parallel sampling the strategist multiple times, we can derive a large amount of aspect-guidance pairs.
@@
Different from the naive wide-horizon thinking that treats the aspects independently, the strategist additionally selects and aggregates the aspects to route the best planning blueprint. From the experiments in Appendix D.2, we find that if we directly increase the number of aspects by sampling the strategist more times, the planner only benefits from considering 3 ~5 aspects without further inference-time scaling capability. It is because more aspects introduce more information but also more noise. To address this, the strategist aggregates the number of aspects into a smaller number and constructs a coherent planning blueprint. As shown in Figure 2, subsequent aspect guidance is determined through the influence of preceding multiple aspects. The routing process shifts the burden of considering numerous aspects from the planner to the strategist, thus enabling the scalability of considering more aspects to further improve the wide-horizon thinking.
@@ 3.2.2 Aspect-Aware Thinking
Based on the planning blueprint constructed by the strategist, the planner sequentially conducts thinking over the aspects in a coherent multi-turn dialogue. With aspect-specific guidance, the planner can conduct a more focused and in-depth analysis over the long context from this aspect. After multiple turns of profound analysis, the planner produces the final plan based on the previous wide-horizon thoughts in the last turn of the conversation.
@@ 3.3 MAoP Distillation for One-Step Wide-Horizon Thinking
Implementing MAoP is complex, involving two models and a multi-turn planning process. We accelerate and simplify this process by distillation. To create high-quality training data for distillation, we employ a powerful teacher model, composed of a strategist and a planner, to generate MAoP samples. From these generated planning trajectories, we extract the strategist’s aspect-specific guidance and then compress the entire aspect-aware thinking and the final aggregation into a single, consolidated output. By finetuning on this distilled data, the model learns to execute complex MAoP planning in a single inference step. This capability is what we term one-step wide-horizon thinking.
4 Causal Evaluation via Agent-based Simulation
As shown in Table 2, previous benchmarks typically rely on static, rule-based metrics, such as pass rates for individual constraints. However, this approach overlooks a fundamental truth: travel is a dynamic, causal process, not a static checklist. Each event, from a delayed train to physical exhaustion on the first day, directly impacts the feasibility and enjoyment of the rest of the journey. By neglecting these critical causal dependencies, existing benchmarks fail to adequately evaluate a plan’s real-world viability and its capacity to meet a traveler’s evolving personal needs.
We introduce Travel-Sim, a novel benchmark framework utilizing agent-based simulation. In this framework, a traveler agent, embodied by the advanced Gemini 2.5-Pro-Exp-0325 google2024gemini , simulates a journey according to a given travel plan. Throughout the simulation, the agent provides continuous, experience-driven feedback. This dynamic, simulation-based approach inherently resolves the causal dependency issues of static benchmarks while capturing personalized, multi-granularity evaluations from an authentic traveler’s perspective.